# Model Inversion Attacks Meet Cryptographic Fuzzy Extractors

### v1

@@ 8 Experimental Evaluation: L2FE-Hash
To be practically usable, L2FE-Hash must have reasonably good accuracy, while defending against model inversion. We evaluate L2FE-Hash for 100 pairs of same user images and 100 pairs of different user images from Celeb-A dataset. From Table XII, we observe that L2FE-Hash has above $60\%$ TPR for both target models (feature extractors) with FPR less than $26\%$. Since the TPR $\geq 50\%$ and FPR $\leq 50\%$, one can employ majority voting over multiple independent samples of the user’s face to boost the accuracy overall as needed, as previously mentioned in Section 7 [53].
@@ 9 Related Work
An embedding inversion attack is a canonical example of model inversion attacks that targets model embeddings [7]. A number of model inversion attacks on image recognition ML models have been developed, often targeting image classification outputs [30, 31, 54, 29, 55]. There are also model inversion attacks on face authentication systems that target image embeddings [14, 13]. We have evaluated $3$ prior attacks representative of SOTA attack accuracy.
Gradient inversion attacks [56] require more than just model outputs and do not meet the original definitions of model inversion [7]. They are not in scope of our defenses.
@@
Many defenses against model inversion modify the training process, not post-processing outputs of ML models already trained [15, 16, 17, 18]. Being highly model-specific and optimization-based, it is hard to theoretically characterize their end security guarantees. Their recently reported effectiveness against existing black-box attacks is fairly limited, with attacks re-identifying $\geq 30\%$ of users from reconstructed images from models that have accuracy over $\geq 80\%$ [17, 15]. Importantly, our work points towards defining properties of cryptographic defenses that generically work with ML models for authentication and in an attack-agnostic way.
The privacy risk from model inversion is immediate, which is our primary concern. There is also re-authentication risk, i.e., using recovered face images to impersonate the user. Liveness detectors during face authentication mitigate such risks, but a line of prior works originating with Xu et. al. [57] shows attacks against such mechanisms. These techniques are complementary to mechanisms we study.
@@
Most post-processing protection mechanisms for embedding vectors are some variant of MRP, or use other non-invertible transformations that do not satisfy fuzzy one-wayness (Item 2). For example, BioHashing is MRP followed by a binary quantization step, and it is known to have pre-image attacks [34, 58, 59]. Index-of-Max Hashing performs random projection on fingerprint minutiae and is also vulnerable [35, 60]. In fact, many protection mechanisms that rely on a distance-preserving transformation are vulnerable to pre-image attacks [61, 62], and thus do not satisfy Item 2. There are also schemes that combine MRP with fuzzy commitment for additional security, but they still do not address inversion attacks that recover approximately close inputs [63, 64].
Similarly, fuzzy hashes, or locality-sensitive hashes [65, 66] aim to find linear projections that preserve local information. For example, Apple’s NeuralHash trains a CNN model to learn such mappings for security applications, but it remains vulnerable to adversarial attacks that force hash collisions [3, 67]. Furthermore, CNN models in general are particularly susceptible to model inversion attacks [39].
Other fuzzy extractor schemes have been proposed, such as the LWE-FE [68]. While being similar to our L2FE-Hash, it has a different decoder construction that targets the ($\ell_{0}$ norm) Hamming distance. Moreover, it assumes the input distribution to be uniform over a small interval or LWE-admissible, which is an extremely strong assumption for real face distributions. To our knowledge, no prior work gives a secure FE construction for $\ell_{2}$ distances under assumptions that even approximately models real-world distributions and yet shows reasonable accuracy in practical setups.
@@
Our L2FE-Hash security definition under full-leakage threat model is not to be confused with post-compromise security. The latter addresses the issue of secure encryption of future messages after key leakage, whereas Item 2 addresses the issue of information leakage about past messages after key leakage [69].
Item 2 is also different from forward secrecy, a property of key-exchange protocols ensuring that compromising a long-term key does not reveal past session keys [70]. The fundamental idea that a compromise should not lead to decrypting past messages is similar to L2FE-Hash, but our setting is different from a key-exchange protocol. We additionally assume past encrypted messages, which in our context are protected embeddings, can be compromised.
Honey encryption (HE) addresses a different threat model from ours [71]. It allows for encryption of passwords with low entropy keys such that decryption with incorrect keys yields honey messages—messages that occur as likely as the original message but are hard to guess without a database breach [71, 72]. We consider the event of full database breach. HE makes it difficult for the adversary to distinguish correct from incorrect decryption, thwarting offline message recovery attacks. It can aid in breach detection by alerting when an adversary unknowingly queries the database with the honeyword. Moreover, password encryption schemes generally do not require locality-sensitivity.
@@ 11 Ethical Considerations
All the datasets used in this paper, Casia-Webface [43], Celeb-A [45] and LFW [47], have licenses which allow their use for face verification tasks for research purposes. We do not foresee ethical issues arising from our work, as we have shown both powerful attacks and a strong defense.
@@ A.1 Formal definition
    Security (Fuzzy one-way): Given $(s,y)$ from $\textsc{Gen}(x)$ where $x\stackrel{{\scriptstyle\mathcal{X}}}{{\leftarrow}}X$, it is hard to find $x^{\prime}$ such that $\text{d}(x,x^{\prime})\leq~t$. In other words, for any computationally bounded adversary $\mathcal{A}$ trying to find such $x^{\prime}$, its advantage is negligible, i.e. $\text{Adv}(\mathcal{A})\leq\text{negl}(\lambda)$ (see Eqn 2 for definition of Adv).
@@
The probability in Item 1 is taken over $x^{\prime}\stackrel{{\scriptstyle\mathcal{X}}}{{\leftarrow}}{B}(x,t)$ and $x\stackrel{{\scriptstyle\mathcal{X}}}{{\leftarrow}}X$. The $\delta$ term in Item 1 can be any value in $[0,1]$. It only affects functional correctness, such as the false positives and negatives in the authentication application. On the other hand, “hardness” in Item 2 is cryptographic, i.e., the probability of an adversary violating Item 2 should be a negligible function in $\lambda$. We consider a computationally bounded adversary that runs in $\text{poly}(\lambda)$ time. We formalize the security game of Item 2 in Section A.2. In Item 3, HILL entropy is the computational notion of min entropy [73].
@@ Appendix B Theoretical Security of L2FE-Hash
We prove that L2FE-Hash satisfies FE security under the assumption that the input distribution is uniform over a bounded support consisting of multiple disjoint $\epsilon$-radius balls, as in Theorem 1. The center of each $\epsilon$-ball can be seen as the representative embedding of a user. Then the $\epsilon$-ball around the center represents the cluster of embeddings corresponding to that user. L2FE-Hash is a FE.
@@ B.1 Definitions
A fuzzy extractor (FE) extracts a uniformly random string with error tolerance. It was first proposed by [25], and [50] relaxed the definition to account for $\delta$ error in correctness. We adapt the definition of FE with $\delta$ errors by [50] for a specific input distribution in Definition 6. Let $M$ be a metric space with distance metric d, and $\mathcal{X}$ be a distribution over $M$.
@@
The strong extractor primitive (Definition 7) is commonly used to construct fuzzy extractors. It can be implemented as a universal hash function, and can extract $\kappa\leq\gamma-2\log(1/\epsilon_{\text{FE}})+2$ many bits of randomness, effectively transforming a high entropy input (Definition 11) into a uniformly distributed output (see Lemma 2.4 in [25]).
Definition 7 (Strong extractor [25, Section 2.5]).
Let $\text{Ext}\colon\{0,1\}^{l}\to\{0,1\}^{\kappa}$ be a polynomial time probabilistic function using $\lambda$ bits of randomness. Ext is an efficient average-case $(l,\gamma,\kappa,\epsilon_{\text{FE}})$-strong extractor if for all pairs of random variables ($X,I$) such that $X$ is an $l$-bit string satisfying $\tilde{H}_{\infty}(X|I)\geq\gamma$, we have $\textbf{SD}((Ext(X;K),K,I)$, $(\mathcal{U}(\{0,1\}^{\kappa}),K,I))\leq\epsilon_{\text{FE}}$, where $K$ is uniform on $\{0,1\}^{\lambda}$.
@@
Definition 8 (R-bounded).
For an integer $R>0$, an $m$-dimensional set $X$ is $R$-bounded if in each dimension, its values are bounded between [R,R][-R,R], i.e. $X\subseteq[-R,R]^{m}$.
@@
Definition 11 (Min Entropy [25, Section 2.4]).
@@ B.2 Proof of Intermediate Theorem
By Minkowski’s Convex Body Theorem, $\frac{\text{vol}(S)}{\det(\mathcal{L})}$ estimates the number of lattice points in set $S$ [74]. Let $S$ be $B_{m}(R)$, the $m$-dim. sphere of radius $R$ inscribed within the $R,R]^{m}$. Then for a random lattice $\mathcal{L}(\mathbf{A})$ where $\mathbf{A}$ is full rank, the number of lattice points in $B_{m}(R)$ is
@@
The third equality holds because $\det(\mathcal{L}(\mathbf{A}))=q^{m-l}$ for a full-rank matrix $\mathbf{A}$, and the last $\approx$ is by Stirling’s approximation. By Lemma 4.4 of [68], $\mathbf{A}$ is full rank with probability at least $1-1/q^{m-l}$ as needed in the Theorem.
@@
The proof is similar to that of Lemma 4 in [52], since $\mathcal{X}$ is also a uniform distribution over its support.
@@ B.3 Proof of L2FE-Hash Security
Since Ext can extract $\kappa\leq\gamma-2\log(1/\epsilon_{\text{FE}})+2$ bits of randomness by Lemma 2.4 of [25], the proof completes. ∎
@@ Appendix C FE Implements the Ideal Primitive
We formally prove Theorem 2 in this section, that a computationally secure FE achieves all the properties required by an ideal primitive. The formal definition of the entropy notion in the utility property is given in Definition 13.
Definition 13 (HILL Entropy [73]).
@@
Utility proof. The FE security property implies a utility argument that translates to the utility property Item 3 of the ideal primitive.
Since $H_{\infty}(\mathcal{U}(\{0,1\}^{\kappa}),\mathcal{P})=H_{\infty}(\mathcal{U}(\{0,1\}^{\kappa}))=\kappa$, and $(\mathcal{R},\mathcal{P})\approx_{\epsilon_{\text{FE}}}(\mathcal{U}(\{0,1\}^{\kappa}),P)$ by FE security, then $H^{\text{HILL}}(\mathcal{R},P)\geq\kappa$ by Definition 13.
@@ Appendix D Prior Attacks Details
Bob attack uses an ID loss function which requires information about the feature extractor, hence making the attack whitebox [14]. Bob’s inversion model uses special blocks in its architecture called DSCasConv which are cascaded convolutions with skip connections.
@@
For GMI and KED-MI attacks, we use the same attack pipeline as [15]. For GMI, we use their pretrained GAN model, and for KED-MI, we train our own GAN model for each target model. In particular, 1000 users were chosen to train the classification layer, with around 27 images per user in the train set, and 3 images per user in the test set. 30,000 images of 6453 users different from those 1000 chosen were used to train the KED-MI GAN model. For Bob, we use only one image per user to train the inversion model, for all 9177 users in our Celeb-A train set, without data augmentation.
To customize the attacks for target models with different protection schemes, we use the same model architecture as the facenet-pytorch model [75], and append a final layer that generates protection scheme outputs from the FaceNet embeddings. For example, the final layer returns $Rx||R.\text{flatten()}$ for the MRP protection scheme, $ss$ for Facial-FE, and $Ab+x||A.\text{flatten()}$ for L2FE-Hash.
@@ Appendix E Supplementary Results
We evaluate the quality and similarity of reconstructed images to the original images by computing the Learned Perceptual Image Patch Similarity (LPIPS) metric [9] and the Fréchet inception distance (FID) metric [48]. We claim that LPIPS is no different from Pipe ASR when using a different face recognition model during evaluation than the one used during training. Hence, we also evaluate Pipe ASR across 3 different feature extractor models. All images are resized to $64\times 64$ before evaluation.
@@
FID scores assess the quality of reconstructed images by comparing the distance between the reconstructed and original image distributions. We use the PyTorch-FID library implementation [76]. As seen in Table IX, FID score for Pipe is much lower than that of Bob, indicating that the distribution of images reconstructed using Pipe is of similar quality to that of original images. It is worth noting that the FID score increases significantly for Bob with MRP and L2FE-Hash protection schemes, unlike Pipe. This is because Bob is not able to produce good-quality face images given these embeddings, whereas PIPE still produces human-looking images that are not close to the original images. Hence, FID does not capture the privacy risk.
@@
Since LPIPS relies on closeness of deep embeddings, we believe that Pipe ASR computed using a feature extractor different from the one used by the target model also represents a similarity metric. We use Facenet, ArcFace, and VGG-Face [77] as feature extractors on the Celeb-A dataset to compute ASR.  Table XIII shows that the reconstructed images are still considered $t$-close to the original images and successfully authenticate with ASR $\geq 76\%$ in most cases. Therefore, the Pipe cross-model ASR is indicative of the performance we would expect from any other ML-based similarity metric like LPIPS.

### v2

@@ 8 Experimental Evaluation: L2FE-Hash
To be practically usable, L2FE-Hash must have reasonably good accuracy, while defending against model inversion. We evaluate L2FE-Hash for 100 pairs of same user images and 100 pairs of different user images from Celeb-A dataset. From Table XII, we observe that L2FE-Hash has above $60\%$ TPR for both target models (feature extractors) with FPR less than $26\%$. Since the TPR $\geq 50\%$ and FPR $\leq 50\%$, one can employ majority voting over multiple independent samples of the user’s face to boost the accuracy overall as needed, as previously mentioned in Section 7 [2011:majority-vote].
@@ 9 Related Work
An embedding inversion attack is a canonical example of model inversion attacks that targets model embeddings [2015:model-inversion]. A number of model inversion attacks on image recognition ML models have been developed, often targeting image classification outputs [2021:ked-mi, 2020:gmi, 2021:vmi, 2023:re-thinking, 2023:plgmi]. There are also model inversion attacks on face authentication systems that target image embeddings [2024:bob, 2023:id3pm-diffusion-inversion-attack]. We have evaluated $3$ prior attacks representative of SOTA attack accuracy.
Gradient inversion attacks [2019:gradient-inversion] require more than just model outputs and do not meet the original definitions of model inversion [2015:model-inversion]. They are not in scope of our defenses.
@@
Many defenses against model inversion modify the training process, not post-processing outputs of ML models already trained [2022:bido, 2021:mid, 2024:midre, 2023:defense-gan]. Being highly model-specific and optimization-based, it is hard to theoretically characterize their end security guarantees. Their recently reported effectiveness against existing black-box attacks is fairly limited, with attacks re-identifying $\geq 30\%$ of users from reconstructed images from models that have accuracy over $\geq 80\%$ [2024:midre, 2022:bido]. Importantly, our work points towards defining properties of cryptographic defenses that generically work with ML models for authentication and in an attack-agnostic way.
The privacy risk from model inversion is immediate, which is our primary concern. There is also re-authentication risk, i.e., using recovered face images to impersonate the user. Liveness detectors during face authentication mitigate such risks, but a line of prior works originating with Xu et. al. [2016:defeat-liveness-detect] shows attacks against such mechanisms. These techniques are complementary to mechanisms we study.
@@
Most post-processing protection mechanisms for embedding vectors are some variant of MRP, or use other non-invertible transformations that do not satisfy fuzzy one-wayness (Item 2). For example, BioHashing is MRP followed by a binary quantization step, and it is known to have pre-image attacks [2004:biohashing, 2005:bh-preimage-attack-withR, 2009:bh-preimage-attack-withoutR]. Index-of-Max Hashing performs random projection on fingerprint minutiae and is also vulnerable [2018:iom-hash, 2020:iom-hash-attack]. In fact, many protection mechanisms that rely on a distance-preserving transformation are vulnerable to pre-image attacks [2024:rp-attack, 2019:genetic-preimage-attack], and thus do not satisfy Item 2. There are also schemes that combine MRP with fuzzy commitment for additional security, but they still do not address inversion attacks that recover approximately close inputs [2019:rp-fc, 2021:rp-fc-noninvert].
Similarly, fuzzy hashes, or locality-sensitive hashes [1999:locality-sensitive-hash, 2014:locality-preserve-hash] aim to find linear projections that preserve local information. For example, Apple’s NeuralHash trains a CNN model to learn such mappings for security applications, but it remains vulnerable to adversarial attacks that force hash collisions [2021:neural-hash, 2022:neural-hash-attack]. Furthermore, CNN models in general are particularly susceptible to model inversion attacks [2022:cnn-recontruction-attack].
Other fuzzy extractor schemes have been proposed, such as the LWE-FE [2020:lwe-fe]. While being similar to our L2FE-Hash, it has a different decoder construction that targets the ($\ell_{0}$ norm) Hamming distance. Moreover, it assumes the input distribution to be uniform over a small interval or LWE-admissible, which is an extremely strong assumption for real face distributions. To our knowledge, no prior work gives a secure FE construction for $\ell_{2}$ distances under assumptions that even approximately models real-world distributions and yet shows reasonable accuracy in practical setups.
@@
Our L2FE-Hash security definition under full-leakage threat model is not to be confused with post-compromise security. The latter addresses the issue of secure encryption of future messages after key leakage, whereas Item 2 addresses the issue of information leakage about past messages after key leakage [2016:post-compromise].
Item 2 is also different from forward secrecy, a property of key-exchange protocols ensuring that compromising a long-term key does not reveal past session keys [2019:forward-secrecy]. The fundamental idea that a compromise should not lead to decrypting past messages is similar to L2FE-Hash, but our setting is different from a key-exchange protocol. We additionally assume past encrypted messages, which in our context are protected embeddings, can be compromised.
Honey encryption (HE) addresses a different threat model from ours [2014:honey-enc]. It allows for encryption of passwords with low entropy keys such that decryption with incorrect keys yields honey messages—messages that occur as likely as the original message but are hard to guess without a database breach [2014:honey-enc, 2024:bernoulli-honeywords]. We consider the event of full database breach. HE makes it difficult for the adversary to distinguish correct from incorrect decryption, thwarting offline message recovery attacks. It can aid in breach detection by alerting when an adversary unknowingly queries the database with the honeyword. Moreover, password encryption schemes generally do not require locality-sensitivity.
@@ 11 Ethical Considerations
All the datasets used in this paper, Casia-Webface [2014:casia-webface], Celeb-A [2015:celeba] and LFW [21012:lfw], have licenses which allow their use for face verification tasks for research purposes. We do not foresee ethical issues arising from our work, as we have shown both powerful attacks and a strong defense.
Acknowledgment
We would like to thank Divesh Aggarwal for insightful discussions that helped clarify the proof of Theorem 1. This research is supported by a Singapore Ministry of Education (MOE) Tier 2 grant MOE-T2EP20124-0007.
@@ A.1 Formal definition
    Security (Fuzzy one-way): Given $(s,y)$ from $\textsc{Gen}(x)$ where $x\stackrel{{\scriptstyle\mathcal{X}}}{{\leftarrow}}X$, it is hard to find $x^{\prime}$ such that $\text{d}(x,x^{\prime})\leq~t$. In other words, for any computationally bounded adversary $\mathcal{A}$ trying to find such $x^{\prime}$, its advantage i.e. $\text{Adv}(\mathcal{A})\leq\text{negl}(\lambda)$ (see Eqn 2 for definition of Adv).
@@
The probability in Item 1 is taken over $x^{\prime}\stackrel{{\scriptstyle\mathcal{X}}}{{\leftarrow}}{B}(x,t)$ and $x\stackrel{{\scriptstyle\mathcal{X}}}{{\leftarrow}}X$. The $\delta$ term in Item 1 can be any value in $[0,1]$. It only affects functional correctness, such as the false positives and negatives in the authentication application. On the other hand, “hardness” in Item 2 is cryptographic, i.e., the probability of an adversary violating Item 2 should be a negligible function in $\lambda$. We consider a computationally bounded adversary that runs in $\text{poly}(\lambda)$ time. We formalize the security game of Item 2 in Section A.2. In Item 3, HILL entropy is the computational notion of min entropy [1999:extractor-universal-hash-hill].
@@ Appendix B Theoretical Security of L2FE-Hash
We prove that L2FE-Hash satisfies FE security under the assumption that the input distribution is uniform over a bounded support consisting of multiple disjoint $\epsilon$-radius balls, as in Theorem 1. The center of each $\epsilon$-ball can be seen as the representative embedding of a user. Then the $\epsilon$-ball around the center represents the cluster of embeddings corresponding to that user.
@@ B.1 Definitions
A fuzzy extractor (FE) extracts a uniformly random string with error tolerance. It was first proposed by [2004:fe], and [2020:fe-impossible] relaxed the definition to account for $\delta$ error in correctness. We adapt the definition of FE with $\delta$ errors by [2020:fe-impossible] for a specific input distribution in Definition 6. Let $M$ be a metric space with distance metric d, and $\mathcal{X}$ be a distribution over $M$.
@@
The strong extractor primitive (Definition 7) is commonly used to construct fuzzy extractors. It can be implemented as a universal hash function, and can extract $\kappa\leq\gamma-2\log(1/\epsilon_{\text{FE}})+2$ many bits of randomness, effectively transforming a high entropy input (Definition 10) into a uniformly distributed output (see Lemma 2.4 in [2004:fe]).
Definition 7 (Strong extractor [2004:fe, Section 2.5]).
Let $\text{Ext}\colon\{0,1\}^{l}\to\{0,1\}^{\kappa}$ be a polynomial time probabilistic function using $\lambda$ bits of randomness. Ext is an efficient average-case $(l,\gamma,\kappa,\epsilon_{\text{FE}})$-strong extractor if for all pairs of random variables ($X,I$) such that $X$ is an $l$-bit string satisfying $\tilde{H}_{\infty}(X|I)\geq\gamma$, we have $\textbf{SD}((\text{Ext}(X;K),K,I)$, $(\mathcal{U}(\{0,1\}^{\kappa}),K,I))\leq\epsilon_{\text{FE}}$, where $K$ is uniform on $\{0,1\}^{\lambda}$.
@@
Definition 10 (Min Entropy [2004:fe, Section 2.4]).
@@ B.2 Proof of Intermediate Theorem
By Minkowski’s Convex Body Theorem, $\frac{\text{vol}(S)}{\det(\mathcal{L})}$ estimates the number of lattice points in set $S$ [1996:minkowski-textbook]. Let $S$ be $B_{m}(R)$, the $m$-dim. sphere of radius $R$ inscribed within the $R,R]^{m}$. Then for a random lattice $\mathcal{L}(\mathbf{A})$ where $\mathbf{A}$ is full rank, the number of lattice points in $B_{m}(R)$ is
@@
The third equality holds because $\det(\mathcal{L}(\mathbf{A}))=q^{m-l}$ for a full-rank matrix $\mathbf{A}$, and the last $\approx$ is by Stirling’s approximation. By Lemma 4.4 of [2020:lwe-fe], $\mathbf{A}$ is full rank with probability at least $1-1/q^{m-l}$ as needed in the Theorem.
@@
The proof is similar to that of Lemma 4 in [2013:dmq], since $\mathcal{X}$ is also a uniform distribution over its support.
@@ B.3 Proof of L2FE-Hash Security
Since Ext can extract $\kappa\leq\gamma-2\log(1/\epsilon_{\text{FE}})+2$ bits of randomness by Lemma 2.4 of [2004:fe], the proof completes. ∎
@@ Appendix C FE Implements the Ideal Primitive
[…]
Definition 12 (HILL Entropy [1999:extractor-universal-hash-hill]).
@@
Utility proof. The FE security property implies a utility argument that translates to the utility property Item 3 of the ideal primitive. Since $H_{\infty}(\mathcal{U}(\{0,1\}^{\kappa}),\mathcal{P})=H_{\infty}(\mathcal{U}(\{0,1\}^{\kappa}))=\kappa$, and $(\mathcal{R},\mathcal{P})\approx_{\epsilon_{\text{FE}}}(\mathcal{U}(\{0,1\}^{\kappa}),P)$ by FE security, then $H^{\text{HILL}}(\mathcal{R},P)\geq\kappa$ by Definition 12.
@@ Appendix D Prior Attacks Details
Bob attack uses an ID loss function which requires information about the feature extractor, hence making the attack whitebox [2024:bob]. Bob’s inversion model uses special blocks in its architecture called DSCasConv which are cascaded convolutions with skip connections.
@@
For GMI and KED-MI attacks, we use the same attack pipeline as [2022:bido]. For GMI, we use their pretrained GAN model, and for KED-MI, we train our own GAN model for each target model. In particular, 1000 users were chosen to train the classification layer, with around 27 images per user in the train set, and 3 images per user in the test set. 30,000 images of 6453 users different from those 1000 chosen were used to train the KED-MI GAN model. For Bob, we use only one image per user to train the inversion model, for all 9177 users in our Celeb-A train set, without data augmentation.
To customize the attacks for target models with different protection schemes, we use the same model architecture as the facenet-pytorch model [2019:facenet-pytorch], and append a final layer that generates protection scheme outputs from the FaceNet embeddings. For example, the final layer returns $Rx||R.\text{flatten()}$ for the MRP protection scheme, $ss$ for Facial-FE, and $Ab+x||A.\text{flatten()}$ for L2FE-Hash.
@@ Appendix E Supplementary Results
We evaluate the quality and similarity of reconstructed images to the original images by computing the Learned Perceptual Image Patch Similarity (LPIPS) metric [2018:LPIPS] and the Fréchet inception distance (FID) metric [2017:fid]. We claim that LPIPS is no different from Pipe ASR when using a different face recognition model during evaluation than the one used during training. Hence, we also evaluate Pipe ASR across 3 different feature extractor models. All images are resized to $64\times 64$ before evaluation.
@@
FID scores assess the quality of reconstructed images by comparing the distance between the reconstructed and original image distributions. We use the PyTorch-FID library implementation [2020:fid-implementation]. As seen in Table IX, FID score for Pipe is much lower than that of Bob, indicating that the distribution of images reconstructed using Pipe is of similar quality to that of original images. It is worth noting that the FID score increases significantly for Bob with MRP and L2FE-Hash protection schemes, unlike Pipe. This is because Bob is not able to produce good-quality face images given these embeddings, whereas PIPE still produces human-looking images that are not close to the original images. Hence, FID does not capture the privacy risk.
@@
Since LPIPS relies on closeness of deep embeddings, we believe that Pipe ASR computed using a feature extractor different from the one used by the target model also represents a similarity metric. We use Facenet, ArcFace, and VGG-Face [2015:vgg-face] as feature extractors on the Celeb-A dataset to compute ASR.  Table XIII shows that the reconstructed images are still considered $t$-close to the original images and successfully authenticate with ASR $\geq 76\%$ in most cases. Therefore, the Pipe cross-model ASR is indicative of the performance we would expect from any other ML-based similarity metric like LPIPS.
Appendix F Reconstructed images
In this section we present some reconstructed images obtained by Bob and Pipe attack on the Celeb-A test dataset. We present reconstructed images for all setups– unprotected embeddings, FacialFE protected embeddings, MRP protected embeddings, and L2FE-Hash protected embeddings. We show the results for Facenet embeddings in Figure 4 and ArcFace embeddings in Figure 5. We observe that both Pipe and Bob can invert unprotected and FacialFE protected embeddings, though the images generated by Bob do not appear to be realistic. This is also reflected by the high FID score for Bob in Section 6. Unlike Bob, Pipe is also able to recover images somewhat similar to the original images from MRP protected embeddings. However, both Pipe and Bob on L2FE-Hash do not recover images similar to the original, demonstrating its security.
Figure 4: Images generated by attacking unprotected, FacialFE protected, MRP protected and L2FE-Hash protected Facenet embeddings. (a) Inversion using Pipe (b) Inversion using Bob. The first column in each subfigure shows the original image.
Figure 5: Images generated by attacking unprotected, FacialFE protected, MRP protected and L2FE-Hash protected ArcFace embeddings. (a) Inversion using Pipe (b) Inversion using Bob. The first column in each subfigure shows the original image.